OpenStudy (anonymous):

I will give a medal.. Find the two independent solution of d^2y/dx^2 + x.dy/dx + y=0 when the series expansion is about the origin

OpenStudy (perl):

suppose the solution is $$ \Large y(x) = a_0 + a_1x + a_2x^2 + a_3 x^3 + ... \\ \text{we want to solve } \\ \Large y ' ' + x y' + y = 0 $$

OpenStudy (perl):

$$ \large y(x) = a_0 + a_1x + a_2x^2 + a_3 x^3 + a_4x^4... \\ \large y'(x) = a_1 + 2a_2x + 3a_3x^2 + 4a_3x^3... \\ \large y ' '(x) = 2a_2 +6a_3x + 12a_4x^2...\\ \large \text{plug these solutions into } \\ \large y ' ' + x y' + y = 0 \large \\ \large \text {we have} \\ ( 2a_2 +6a_3x + 12a_4x^2... )+ x(a_1 + 2a_2x + 3a_3x^2 + 4a_3x^3... ) \\ + ( a_0 + a_1x + a_2x^2 + a_3 x^3 + a_4x^4...)=0 $$

OpenStudy (perl):

$$( 2a_2 +6a_3x + 12a_4x^2... )+ x(a_1 + 2a_2x + 3a_3x^2 + 4a_3x^3... ) \\ + ( a_0 + a_1x + a_2x^2 + a_3 x^3 + a_4x^4...)=0\\ \iff \\ (2a_2 + a_0) + (6a_3 + a_1+ a_1)x + (12a_4 +2a_2 + a_2)x^2 + ...=0 \\\iff\\ (2a_2 + a_0) + (6a_3 +2a_1)x + (12a_4 +3a_2)x^2 + ...=0 \\ \iff \\ \text {since the right side is equal to } \\ 0 + 0x +0x^2 + 0x^3 + ...\\ \text {equating coefficients we have}\\ 2a_2 + a_0 =0\\ 6a_3 +2a_1 = 0\\ 12a_4 + 3a_2 = 0 $$

OpenStudy (perl):

it is easier if we use sum notation

OpenStudy (anonymous):

Thanks prof. Am learning more. Please ride on sir

OpenStudy (anonymous):

Ok sir.Sir from the first method you used, what are then the independent solutions

OpenStudy (perl):

$$ \text{we want to solve } \\ \Large y ' ' + x y' + y = 0 \\ \text {Assume your solution is } \\ \Large y(x)=\sum_{n=0}^{\infty}a_n x^n\\ \text {then it follows that} \\ \Large y'(x) = \sum_{n=1}^{\infty}n \cdot a_n x^{n-1} \\ \Large y''(x) = \sum_{n=2}^{\infty}n(n-1) \cdot a_n x^{n-2}\\ \text {Substitute: }\Large y ' ' + x y' + y = 0 \\ \large \sum_{n=2}^{\infty}n(n-2) \cdot a_n x^{n-2} + x\sum_{n=1}^{\infty}n \cdot a_n x^{n-1} + \sum_{n=0}^{\infty}a_n x^n=0\\ \large \iff \\ \large \sum_{n=2}^{\infty}n(n-2) \cdot a_n x^{n-2} + \sum_{n=1}^{\infty}n \cdot a_n x^{n} + \sum_{n=0}^{\infty}a_n x^n=0\\ \large \iff \\ \large \sum_{n=0}^{\infty}(n+2)(n+1) \cdot a_{n+2} x^{n} + \sum_{n=1}^{\infty}n \cdot a_n x^{n} + \sum_{n=0}^{\infty}a_{n+2} x^n=0\\ \large \iff\\ \large \sum_{n=0}^{\infty}(n+2)(n+1) \cdot a_{n+2} x^{n} + \sum_{n=1}^{\infty}n \cdot a_n x^{n} + \sum_{n=0}^{\infty}a_n x^n=0\\ $$

OpenStudy (anonymous):

Ok.i understand sir but I instead of using

OpenStudy (perl):

make sure there are no errors, we need this for the next step

OpenStudy (anonymous):

using the sum notationn can you deduce the two linear independent soluctions from the first method

OpenStudy (perl):

$$ \iff \\ \ (2)(1)a_2x^0 + \sum_{n=1}^{\infty}(n+2)(n+1)a_{n+2}x^n + \sum_{n=1}^{\infty}n \cdot a_n x^{n} + a_0 x^0 + \sum_{n=1}^{\infty}a_n x^n=0 \\ \iff \\ \large (2a_2 +a_0) + \sum_{n=1}^{\infty}[(n+2)(n+1)a_{n+2} + n\cdot a_n + a_n] x^n =0\\ \iff \\ \large (2a_2 +a_0) + \sum_{n=1}^{\infty}[(n+2)(n+1)a_{n+2} + (n+1)\cdot a_n] x^n =0 \\ \large \text{Equating coefficients this gives us: } \\ \large (2a_2+a_0)= 0\\ \large (n+2)(n+1)a_{n+2} + (n+1)\cdot a_n=0 \\ \text{Therefore}\\ \large a_0 = -2a_2 \\ \large a_{n+2} = \frac{-a_n(n+1)}{(n+1)(n+2)} = \frac{-a_n}{n+2} \\ $$

OpenStudy (anonymous):

is that the answer sir?

OpenStudy (perl):

made a typo

OpenStudy (perl):

$$ \text {look at the even terms }\\ \Large a_2 = \frac{-a_0}{2} \\ \Large a_4 = \frac{-a_2}{(2+2)}= \frac{-(\frac{-a_0}{2})}{4}= \frac {a_0}{8} \\ \Large a_6 = \frac{-a_4}{(4+2)}= \frac{-(\frac{a_0}{8})}{6}= -\frac {a_0}{48} \\ ... \\ \Large a_{2n}= \frac{a_0}{2^n~n!}(-1)^n \\ ~~~\\ \text {look at odd terms }\\ \Large a_3 = \frac{-a_1}{1+2}= \frac{-a_1}{3}\\ \Large a_5 = \frac{-a_3}{3+2}= \frac{-(\frac{-a_1}{3})}{5}= \frac {a_1}{15} \\ \Large a_6 = \frac{-a_4}{5+2}= \frac{-(\frac{a_1}{15})}{7}= -\frac {a_1}{105} \\ ... \\ \Large a_{2n+1}=\frac{a_1(-1)^n (2n+1)!}{2^n n!} $$

OpenStudy (anonymous):

OK SIR BUT CAN I ASK ANOTHER QUESTION HERE?

OpenStudy (perl):

better put a new post

OpenStudy (anonymous):

BUT PLEASE HELP ME ON THAT ...... THANKS NEW POST COMING

OpenStudy (anonymous):

to get y1 , let a_0 = a_1 = 1 to get y2 , let a_0 = -a_1 = 1 PLEASE COMPLETE IT SIR

OpenStudy (perl):

$$\large \text{plug in the a_n that we found}\\ \large y(x)=\sum_{n=0}^{\infty}a_n x^n\\ \large = a_0 + a_1x + a_2x^2 + a_3x^3 + a_4x^4...\\ \large = a_0 + a_1x + \frac{-a_0}{2}x^2 + \frac{-a1}{3}x^3 + \frac{a_0}{8}x^4 +\frac{a_0}{15}x^4... = $$

OpenStudy (perl):

now we can factor

OpenStudy (anonymous):

SIR CAN YOU WORK ON THIS FIND THE FIRST FIVE TERM IN THE SOLUCTION DY/DX + (1+X^2)Y= SINX WITH Y(0)= a

OpenStudy (anonymous):

?

OpenStudy (perl):

notice that there are two independent constants there $$ a_0, a_1 $$

OpenStudy (perl):

$$ \large \text{plug in the a_n that we found}\\ \large y(x)=\sum_{n=0}^{\infty}a_n x^n\\ \large = a_0 + a_1x + a_2x^2 + a_3x^3 + a_4x^4...\\ \large = a_0 + a_1x + \frac{-a_0}{2}x^2 + \frac{-a1}{3}x^3 + \frac{a_0}{8}x^4 +\frac{a_0}{15}x^4...\\ = a_0 + a_1x + \frac{-a_0}{2}x^2 + ... +\frac{a_0(-1)^n}{2^n n!} +\frac{a_0(-1)^{n+1}(2n+1)!}{2^n n!}... \\ = a_0 \Big\{1 - \frac{x^2}{2} +\frac{x^4}{8} - ... \Big\} + a_1\Big\{x - \frac{x^3}{3} +\frac{x^5}{15} - ... \Big\} \\ \large = a_0 \sum_{n=0}^{\infty}\frac{(-1)^nx^{2n}}{2^n n!}+ a_1 \sum_{n=0}^{\infty}\frac{(-1)^nx^{2n+1}(2n+1)!}{2^n n!} $$

OpenStudy (anonymous):

so sir what will be my conclusion

OpenStudy (perl):

if you set a_0 = 0, and a_1 = 1, you will have one solution if you set a_0 = 1 and a_0 = 0 you will have another solution

OpenStudy (anonymous):

ok sir

OpenStudy (anonymous):

thank you sir

OpenStudy (perl):

$$ \large \text{If we substitute in the} ~a_n \text{that we found recursively above}\\ \large y(x)=\sum_{n=0}^{\infty}a_n x^n\\ \large = a_0 + a_1x + a_2x^2 + a_3x^3 + a_4x^4...\\ \large = a_0 + a_1x + \frac{-a_0}{2}x^2 + \frac{-a1}{3}x^3 + \frac{a_0}{8}x^4 +\frac{a_0}{15}x^4...\\ = a_0 + a_1x + \frac{-a_0}{2}x^2 + ... +\frac{a_0(-1)^n}{2^n n!} +\frac{a_0(-1)^{n+1}(2n+1)!}{2^n n!}... \\ = a_0 \Big\{1 - \frac{x^2}{2} +\frac{x^4}{8} - ... \Big\} + a_1\Big\{x - \frac{x^3}{3} +\frac{x^5}{15} - ... \Big\} \\ \large = a_0 \sum_{n=0}^{\infty}\frac{(-1)^nx^{2n}}{2^n n!}+ a_1 \sum_{n=0}^{\infty}\frac{(-1)^nx^{2n+1}(2n+1)!}{2^n n!} \\ \text {Therefore} \\ \large y(x) =a_0 \sum_{n=0}^{\infty}\frac{(-1)^nx^{2n}}{2^n n!}+ a_1 \sum_{n=0}^{\infty}\frac{(-1)^nx^{2n+1}(2n+1)!}{2^n n!} \\ \large \text{if } a_0 = 1 , a_1 = 0 \\\Large y_1(x)= \sum_{n=0}^{\infty}\frac{(-1)^nx^{2n}}{2^n n!} \\\Large\text {if } a_0= 0 ,a_1 = 1 \\\Large y_2(x)=\sum_{n=0}^{\infty}\frac{(-1)^nx^{2n+1}(2n+1)!}{2^n n!} $$

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

can you start the other now or do you need some rest sir

OpenStudy (anonymous):

the nth coefficient of a series is given by A_n= (1.5.9.13....(4n+1))/2^n express a_n in terms of the gamma funtion

OpenStudy (perl):

interesting, the denominator for a_2n+1 is (2n+1)!! , double factorial

OpenStudy (anonymous):

dont understant that sir

OpenStudy (perl):

the coefficient 3,15,105 we found above are product of odd integers 3 = 3 * 1 15 = 5 * 3 * 1 105 = 7 * 5 * 3 * 1 ...

OpenStudy (anonymous):

ok but do not understand what next you are ding

OpenStudy (anonymous):

is the soluction for Y2(x) not correct?

OpenStudy (anonymous):

ok

OpenStudy (perl):

copy and paste that link

OpenStudy (anonymous):

does it mean we are ccorrect sir

OpenStudy (perl):

yes

OpenStudy (anonymous):

sir can you please work on any of the two last questions ?

OpenStudy (perl):

@SithsandGiggles @Michele_Laino if you can look over, i would greatly appreciate (for errors)

OpenStudy (perl):

there is a nice simplification for y_1(x) and y_2(x)

OpenStudy (perl):

$$ \\\Large y_1(x)= \sum_{n=0}^{\infty}\frac{(-1)^nx^{2n}}{2^n n!} \\ \Large = \sum_{n=0}^{\infty}\frac{(\frac{-x^{2}}{2})^n}{n!}\\ \LARGE =e^{\frac{-x^2}{2}} \\[0.5 in] \\\Large y_2(x)=\sum_{n=0}^{\infty}\frac{(-1)^nx^{2n+1}(2n+1)!}{2^n n!} \\ \text{this one doesn't have a nice simplification} $$

OpenStudy (perl):

Again to recap I got for the general solution: $$ \\\Large y(x)=a_0e^{\frac{-x^2}{2}} + a_1\sum_{n=0}^{\infty}\frac{(-1)^nx^{2n+1}(2n+1)!}{2^n n!} $$

OpenStudy (perl):

if you can find a closed form or some kind of interesting approach for the latter term, maybe in terms of gamma function.

OpenStudy (anonymous):

@perl I think there's just a slight mistake, where \(y''=\sum\limits_{n=2}^\infty n(n-1)a_nx^{n-2}\), you had substituted \(\sum\limits_{n=2}^\infty n\color{red}{(n-2)}a_nx^{n-2}\).

OpenStudy (perl):

@SithsAndGiggles yes i see. I don't understand how these typos happen. and i can't edit it now. i need a large latex editor so i can preview this. this was quite a tedious task using Latex. by the way, how come youre name doesn't show up when i type @ , hmm, well my name doesn't appear either

OpenStudy (perl):

Did you see a simple expression for y_2(x)? im going to check wolfram and compare their approximation to solving the linear diff. y ' ' + x y ' + y = 0 The subject of power series is quite vast though, we are barely scratching the surface here, i imagine/

OpenStudy (anonymous):

ok sir thanks.

OpenStudy (anonymous):

but this is an exam question. where can i stop from the solution given, and i understood very well how you solved it but can you explain why ∑n=2∞n(n−2)⋅anxn−2 became ∑n=0∞(n+2)(n+1)⋅an+2xn

OpenStudy (mendicant_bias):

[SIR INTENSIFIES]

OpenStudy (anonymous):

hello bendicent can you help me with my open question ?

OpenStudy (anonymous):

@perl if you're looking for a simpler expression for the answer, you can determine that by solving the ODE a different way. \[\frac{d^2y}{dx^2}+x\frac{dy}{dx}+y=0~~\iff~~\frac{d}{dx}\left[\frac{dy}{dx}+xy\right]=0\] which becomes linear in \(y\) upon integrating: \[\frac{dy}{dx}+xy=C_1\] Finding the integrating factor: \[\ln\mu(x)=\int x\,dx~~\implies~~\mu(x)=e^{x^2/2}\] hence \[e^{x^2/2}\frac{dy}{dx}+xe^{x^2/2}y=C_1e^{x^2/2}~~\iff~~\frac{d}{dx}\left[e^{x^2/2}y\right]=C_1e^{x^2/2}\] Integrating again and solving for \(y\) yields \[\newcommand{erf}{\,\text{erf}\,}y=C_1e^{-x^2/2}\int e^{x^2/2}\,dx+C_2e^{-x^2/2}\] where \[\int e^{x^2/2}\,dx=\sqrt\frac{\pi}{2}\erf\frac{x}{\sqrt 2}=\sum_{n=0}^\infty \frac{\left(-\frac{1}{2}\right)^nx^{2n+1}}{n!(2n+1)}\]